OpenStudy (anonymous):

Find the indicated output for this function. if f(x)=x^2+2x+3, find f(-2) @katlin95 @Sceenie @mathstudent55

OpenStudy (anonymous):

\[f(x)=x^2+2x+3\]\[f(\diamondsuit)=\diamondsuit^2+2\times \diamondsuit+3\]\[f(-2)=(-2)^2+2\times (-2)+3\]

OpenStudy (anonymous):

This is like the 5th one ive done , sub -2 for x . then solve

OpenStudy (anonymous):

how is that @Sceenie